### v2

# AC-LIO: Towards Asymptotic and Consistent Convergence in LiDAR-Inertial Odometry

Abstract
Existing LiDAR-Inertial Odometry (LIO) methods typically utilize the prior state trajectory derived from the IMU integration to compensate for the motion distortion within LiDAR frames. However, discrepancies between the prior and actual trajectory can lead to residual distortions that compromise the consistency of the LiDAR frame with its corresponding geometric environment. This imbalance may result in pointcloud registration becoming trapped in local optima, thereby exacerbating drift during long-term and large-scale localization. To address the issue, we propose a novel asymptotically and consistently converging LIO framework dubbed AC-LIO. Our key idea is to back propagate current update term based on the prior state chain, and asymptotically compensate for the residual distortion during iteration. Moreover, considering the weak correlation between previous error and current distortion, we establish convergence criteria based on the pointcloud constraints to regulate the backpropagation. This method of guiding asymptotic distortion compensation using convergence criteria subtly enhances the consistent convergence of pointcloud registration, futher improving the accuracy and robustness of LIO system. Extensive experiments demonstrate that our AC-LIO framework significantly promotes consistent convergence in state estimation compared to prior arts, with about 30.4% reduction in average RMSE over the second best result, leading to marked improvements in the accuracy of long-term and large-scale localization and mapping.
I INTRODUCTION
With the advancements in Simultaneous Localization and Mapping (SLAM) technology, robots are increasingly achieving precise perception of their external environments and their own motion[1, 2, 3]. As a vital subset of SLAM, LiDAR-Inertial Odometry (LIO) has gained widespread application in autonomous driving [4], unmanned logistics [5], humanoid robot [6] and other fields, for its high robustness and accuracy [7, 8]. Compared to Visual-Inertial Odometry, the LIO system can directly obtain depth observations of surroundings [9], without the need to estimate the depth of feature points [10, 11], and exhibits better adaptability to different lighting conditions [12, 13].
Figure 1: (a) A principal comparison of our AC-LIO with traditional approaches. Conventional LIO typically considers an initial distortion compensation via IMU, yet the residual distortion may prevent the registration from further convergence. In contrast, our AC-LIO employs an asymptotic backpropagation to promote consistent convergence in state estimation. (b) Large scale localization and mapping result of BotanicGarden dataset. We achieve better accuracy performance than other benchmark frameworks on its sequence 1008_01 with VLP16 LiDAR.
However, as LIO system continues to evolve, there still remain several fundamental issues to be addressed [14, 15, 16]. One of them is the conflict between continuous motion and rigid pointcloud registration [17]. Since the acquisition of LiDAR frames is often accompanied with the ego-motion of the vehicle, the LiDAR points with different timestamps actually lie in different coordinate systems.
As shown in Fig. 1, traditional methods [18, 19, 20] typically only consider using IMU integration for the initial compensation of motion distortion, without further optimization of the residual distortion in the state estimation. This residual distortion will compromise the consistency between the LiDAR frame with its corresponding geometric environment, and it grows with increasing motion non-linearity. Therefore, in conventional LIO systems, such residual error may suppress further convergence of pointcloud registration during the iterative update, ultimately impacting the odometry accuracy and causing severe drift.
In this letter, we propose AC-LIO, a novel asymptotically and consistently converging LIO framework. Our AC-LIO enjoys two key technical contributions. First, during the iterative state estimation, we introduce an asymptotic distortion compensation (Sec. III-C) method that backwards propagate the current update term based on the prior state chain, and asymptotically compensate the residual distortion before next iteration . This approach could enhance the consistency between the current LiDAR frame and the represented geometric environment, and avoid the pointcloud registration from falling into local optimum due to the residual distortion, thereby promoting the consistent convergence of pointcloud registration and improving the accuracy of long-term and large-scene localization.
Second, considering the weak correlation between initial error and motion distortion of current frame, we also propose the convergence criteria (Sec. III-D) based on the point-to-plane residuals to control the back propagation. The criteria is used to assess the convergence of the pointcloud registration and ensure that the backpropagation and asymptotic compensation only happens under the condition of sufficient convergence in previous LiDAR frame, hence guaranteeing the robustness of LIO system. Compared to other strategies such as continuous-time estimation or frame segmentation, our method leverages this criteria to construct a feedback mechanism that promotes state convergence without additional estimation, achieving a balance between localization accuracy and computational efficiency (Sec. IV-C). It can be seamlessly embedded into the iterative Kalman Filter framework [21], offering great flexibility and scalability.
In summary, our main contributions are as follows:
  • We propose a novel framework AC-LIO with asymptotic distortion compensation that showing plausible consistent convergence in long-term and large-scale scenarios.
  • We introduce the convergence criteria based on point-to-plane residuals to evaluate the convergence and guide back propagation. Back propagation is only performed upon sufficient convergence of previous registration, thus further enhancing the system robustness.
  • Compared to other optimization strategies for motion distortion, our framework introduce no extra estimates or states, making the asymptotic compensation less computationally expensive but highly flexible.
  • We conduct extensive experiments on AC-LIO in a series of real-world datasets with various scenarios. The results show that our method achieves higher accuracy compared to current state-of-the-art frameworks.
II RELATED WORK
LiDAR-Only Odometry. Studies such as LOAM [22] and F-LOAM [23] has achieved significant performance with the assumption of constant velocity motion during the corresponding time of LiDAR frames. However, this assumption only works effectively when LiDAR frame interval is enough short and the motion is slow and smooth. In case of drastic velocity variation, the constant velocity model cannot accurately characterize the current motion state, and this will result in increased drift and even odometry divergence. CT-ICP [24] optimize the start and end states of LiDAR frame, and construct a continuous time trajectory by interpolation. But the real movement may not follow the interpolated trajectory, as well as the logical constraint across adjacent frames is hard to determine. I2EKF-LO [25] utilizes extra prediction iteration to estimate motion distortion, yet the update of the process noise still depends on the hyperparameters. Traj-LO [26] leverages the high frequency of LiDAR points and employs more linear segments within frame. This approach allows more accurate trajectory characterization and suppression of motion aberrations during high speed motion, but also introduces more computational cost.
Discrete State LiDAR-Inertial Odometry. For LiDAR-inertial odometry, the IMU observation can provide relatively reliable prior state information. FAST-LIO2 [27] employs the IMU state backpropagation to correct the distorted pointcloud, and then conducts state estimation through Error State Kalman Filter. LIO-SAM [28] utilizes the IMU pre-integration to de-skew pointcloud, and fuses multiple constraints including LiDAR, IMU, and GPS through the factor graph. In the subsequent LIO frameworks, SR-LIO [29] adopts a LiDAR frame reconstruction strategy, while AS-LIO [30] designs an adaptive sliding window guided by spatial overlap degree, both of which can boost the system update frequency and lower the impact of motion distortion. Semi-Elastic-LIO [31] replaces the interpolation model with IMU motion constraints on the basis of CT-ICP framework, thereby improving the stability of the inter-frame logical constraints. Although these methods can mitigate the impact of pointcloud distortion to some extent, there are still some limitations without considering the further optimization of residual motion distortion. Point-LIO [32] adopts a point-by-point update strategy to avoid the motion distortion issue and realize extremely high-frequency odometry. Compared to frame based registration, the point-by-point strategy is more likely to affected by dynamic objects and degenerate in adverse scenarios.
Continuous State LiDAR-Inertial Odometry. To characterize trajectorie more precisely in state estimation, many frameworks come up with new optimization solutions. CLINS [33] introduces the B-spline [34] for continuous time state estimation, while SLICT [35] employs continuous-time factors to represent LiDAR constraints within a sliding window. These frameworks further optimize motion distortion with continuous-time model. However, the unconstrained trajectory would not always conform to the predefined model due to time resolution limitations in high-speed, dynamic motion. Moreover, the continuous-time optimization strategy may introduce a higher computational cost. DLIO [36] uses analytical equations to construct a continuous trajectory for precise motion correction, after coarse deskewing through IMU integration. But the method remains limited by IMU’s inherent accuracy and neglects further undistortion during the registration process.
In contrast, our proposed AC-LIO framework utilizes the convergence criteria to guide the backpropagation process, to achieve the asymptotic compensation of residual motion distortion. As a result, our approach can promote the consistent convergence of pointcloud registration during iterative update, further improving the accuracy over long-term and large-scale scenarios while ensuring the system robustness.
III METHODOLOGY
TABLE I: Important Notations
Symbols Definition
[…]
$I,L,G$ The IMU, LiDAR, and global coordinate frame.
$\mathbf{x}^{gt}$ The ground truth of state $\mathbf{x}$.
$\mathbf{x}_{i}$ The vector of state $\mathbf{x}$ at timestamp $i$.
$\hat{\mathbf{x}}$ The priori estimation of state $\mathbf{x}$.
$\hat{\mathbf{x}}^{j},\delta\hat{\mathbf{x}}^{j}$ The $j$-th update of state $\mathbf{x}$ and error state $\delta\mathbf{x}$.
$\bar{\mathbf{x}},\check{\mathbf{x}}$ The updated state $\mathbf{x}$ with or without back propagation.
$\mathbf{F_{a}}^{i|i+n}$ The propagation matrix from state $i$ to $i+n$.
$\mathbf{G}_{i}^{i+n}$ The backward gain matrix of state $i$ with respect to $i+n$.
$\mathbf{z},z$ The vector and its scalar value of point-to-plane residual.
$\phi$ The incident angle of single laser beam on the fitted plane.
$\sigma_{t}$ The expectation of average point-to-plane residual.
III-A System Overview
Figure 2: (a) The system overview of our AC-LIO framework. During iterative ESKF, asymptotic compensation of pointcloud distortion is conducted based on the convergence criteria, to promote consistent convergence of state estimation. (b) The on-manifold propagation. The blue shows the forward propagation of IMU integration. The green indicates the backward recursion of the update term according to the propagation matrix and noise relation.
The system overview of our AC-LIO framework is shown in Fig. 2, and the important notations is shown in Table I. First, the IMU observations are fed into the forward propagation module to obtain prior trajectories. Then the LiDAR data utilize the prior state chain for initial compensation and then sent to the Error State Kalman Filter (ESKF) for state estimation [37], where the average point-to-plane residual (APR) is calculated after each update.
During the iteration, if the APR of previous LiDAR frame is lower than the convergence criteria and the current APR exceeds it, then the backpropagation of update residual will be executed. The LiDAR frame will be re-compensated to promote the convergence of registration after each backpropagation. Otherwise, if the conditions are not met, the backpropagation will be skipped and iteration continues. Once reaching the maximum of iterations, the system will merge the LiDAR frame under the posterior pose into pointcloud map and publish the current odometry.
III-B ESKF State Estimation
For efficiency and accuracy concerns, the AC-LIO we proposed utilizes the ESKF framework for state estimation. Based on the kinematic model and IMU measurement model, the state $\mathbf{x}$ , input $\mathbf{i}$, and noise $\mathbf{n}$ can be defined as follows:
$\begin{split}\mathbf{x}&\doteq\begin{bmatrix}\mathbf{R}_{I}^{G}&\mathbf{p}_{I}^{G}&\mathbf{v}_{I}&\mathbf{b}_{\bm{\omega}}&\mathbf{b}_{\mathbf{a}}&\mathbf{g}^{G}\end{bmatrix}^{T}\\ \mathbf{i}&\doteq\begin{bmatrix}{\bm{\omega}}_{m}&{\mathbf{a}}_{m}\end{bmatrix}^{T},~{}~{}\mathbf{n}\doteq\begin{bmatrix}\mathbf{n}_{\bm{\omega}}&\mathbf{n}_{\mathbf{a}}&\mathbf{n}_{\mathbf{b}\bm{\omega}}&\mathbf{n}_{\mathbf{b}\mathbf{a}}\end{bmatrix}^{T}\end{split}$ (1)
[…]
$\begin{gathered}\quad\quad\delta\mathbf{x}_{i}\doteq\begin{bmatrix}\delta{\bm{\theta}}&\delta\mathbf{p}_{I}^{G}&\delta\mathbf{v}_{I}&\delta\mathbf{b}_{\bm{\omega}}&\delta\mathbf{b}_{\mathbf{a}}&\delta\mathbf{g}^{G}\end{bmatrix}^{T}\\ \quad\delta\hat{\mathbf{x}}_{i+1}=\mathbf{F_{a}}\delta\check{\mathbf{x}}_{i}+\mathbf{F_{n}}\mathbf{n}_{i}\\ \quad\hat{\mathbf{P}}_{i+1}=\mathbf{F_{a}}\check{\mathbf{P}}_{i}\mathbf{F_{a}}^{T}+\mathbf{F_{n}}\mathbf{Q}\mathbf{F_{n}}^{T}\end{gathered}$ (2)
Assume the extrinsic $\mathbf{T}_{L}^{I}$ between LiDAR and IMU is already known. For a point $\mathbf{p}$ in LiDAR frame, its neighboring cluster can be searched by KNN in pointcloud map. If a plane $S$ can be fitted to this cluster, then we will obtain its normal vector $\mathbf{u}$. Note one point of the cluster as $\mathbf{q}$ and $(\cdot)_{\wedge}$ is the upper triangular matrix of vector, then the point-to-plane residual $\mathbf{z}$ and measurement matrix $\mathbf{H}$ are as follows:
$\begin{gathered}\mathbf{z}=\mathbf{u}(\hat{\mathbf{p}}^{G}-\mathbf{q}^{G})^{T}\\ \mathbf{H}=\begin{bmatrix}-\mathbf{u}\hat{\mathbf{R}}_{I}^{G}(\mathbf{T}_{L}^{I}\mathbf{p}^{L})_{\wedge}&\mathbf{u}&\mathbf{0}_{3\times 12}\end{bmatrix}\end{gathered}$ (3)
Considering the nonlinear errors in the $\mathbf{F_{a}}$, $\mathbf{F_{n}}$ and $\mathbf{H}$, the estimation may require iteration to promote convergence. Let $\mathbf{A}$ denote the Jacobian of update component with respect to $\delta\mathbf{x}$, and according to [18], the update equation is:
$\begin{gathered}\mathbf{K}=\mathbf{P}\mathbf{H}^{T}(\mathbf{H}\mathbf{P}\mathbf{H}^{T}+\mathbf{R})^{-1}\\ \hat{\mathbf{x}}_{i+1}^{j+1}=\hat{\mathbf{x}}_{i+1}^{j}\boxplus[-\mathbf{K}\mathbf{z}-(\mathbf{I}-\mathbf{K}\mathbf{H})(\mathbf{A}^{j})^{-1}(\hat{\mathbf{x}}_{i+1}^{j}\boxminus\hat{\mathbf{x}}_{i+1})]\end{gathered}$ (4)
Once the update term is less than a preset threshold, the solution can be considered to converge.
$\begin{gathered}\check{\mathbf{x}}_{i+1}\Leftarrow\hat{\mathbf{x}}_{i+1}^{j+1},\;\check{\mathbf{P}}_{i+1}\Leftarrow(\mathbf{I}-\mathbf{K}\mathbf{H})\hat{\mathbf{P}}_{i+1}\end{gathered}$ (5)
With the iterative update of ESKF, our AC-LIO could further promote the consistent convergence of the state estimation via the asymptotic distortion compensation and convergence criteria introduced in following sections.
III-C Asymptotic Distortion Compensation
Compared to the noise of LiDAR observations, motion distortion may introduce bias in pointcloud registration, making it difficult to utilize the statistical properties of pointcloud to mitigate the adverse effects on state estimation.
During the on-manifold propagation of state $\mathbf{x}$, the noise of current error state $\delta\hat{\mathbf{x}}_{i+1}$ consists of the propagation of previous state $\delta\hat{\mathbf{x}}_{i}$ and the accumulated process noise. With reference to the Rauch-Tung-Striebel fixed interval smooth strategy [39], we can back propagate the update term $\delta\mathbf{x}$ based on the prior state chain. The contribution of previous state in current update term just corresponds to the proportion of propagation covariance $\mathbf{F_{a}}\check{\mathbf{P}}_{i}\mathbf{F_{a}}^{T}$ within current covariance $\hat{\mathbf{P}}_{i+1}$. Then, the gain matrix $\mathbf{G}$ and the backward recursion of $\delta\mathbf{x}$ can be denoted as follows:
$\begin{gathered}\mathbf{G}=\check{\mathbf{P}}_{i}\mathbf{F_{a}}^{T}\hat{\mathbf{P}}_{i+1}^{-1}\\ \delta\hat{\mathbf{x}}_{i}^{j}=\mathbf{G}\cdot\delta\hat{\mathbf{x}}_{i+1}^{j}\end{gathered}$ (6)
Furthermore, this sequential backpropagation method can be extend to any states within the LiDAR frame. It allows us to selectively update several anchor in the prior state chain, thus reducing the computational complexity of backpropagation and distortion compensation. As for any states $\mathbf{x_{i}}$ and $\mathbf{x_{i+n}}$ in the prior state chain, the propagation matrix and backward gain matrix can be noted as $\mathbf{F_{a}}^{i|i+n}$ and $\mathbf{G}_{i}^{i+n}$, then the backpropagation relationship should be:
$\begin{gathered}\mathbf{F_{a}}^{i|i+n}=\prod_{k=i}^{i+n-1}\mathbf{F_{a}}^{k|k+1}\\ \mathbf{G}_{i}^{i+n}=\check{\mathbf{P}}_{i}(\mathbf{F_{a}}^{i|i+n})^{T}(\hat{\mathbf{P}}_{i+n})^{-1}\\ \check{\mathbf{x}}_{i}^{j}=\check{\mathbf{x}}_{i}\boxplus(\mathbf{G}_{i}^{i+n}\cdot\delta\hat{\mathbf{x}}_{i+n}^{j})\end{gathered}$ (7)
After each backpropagation of the update term $\hat{\mathbf{x}}_{i+n}^{j}$, the residual pointcloud distortion can continue to be compensated based on the smoothed trajectory. For the point $\mathbf{p}^{L}$ in LiDAR frame, and the state corresponding to its timestamp $t$ after the $j$th backpropagation can be denoted as $\hat{\mathbf{x}}_{t}^{j}$, then the updated coordinates can be noted as:
$\begin{gathered}\hat{\mathbf{p}}_{j}^{G}=\mathbf{T}(\hat{\mathbf{x}}_{t}^{j})\mathbf{T}_{L}^{I}\mathbf{p}^{L}\end{gathered}$ (8)
Where $\mathbf{T}(\mathbf{x})$ represents the transformation matrix of state $\mathbf{x}$ in frame $G$. As the iterative ESKF converges, the final solution of the backward propagation can be expressed as:
$\begin{gathered}\bar{\mathbf{x}}_{i}\Leftarrow\check{\mathbf{x}}_{i}^{j+1},\;\bar{\mathbf{P}}_{i}\Leftarrow\mathbf{G}_{i}^{i+n}\check{\mathbf{P}}_{i+n}(\mathbf{G}_{i}^{i+n})^{T}\end{gathered}$ (9)
In practice, if there is no need to reuse the previous state queue in next frame, the covariance update after backpropagation can be skipped for efficiency.
III-D Convergence Criteria
Figure 3: The relationship between original ranging variance of LiDAR point and point-to-plane residual variance, where the residual $z$ is a scalar.
To exploit the statistical properties of pointcloud, during the iterative state estimation, we select the mean value of the point-to-plane residual (abbreviated as APR above), as the criteria for evaluating the convergence of pointcloud registration. For simplification, assuming the pointcloud registration fully converges to the global optimum and only considering the ranging accuracy of LiDAR, we can roughly model the expectation of the APR and provide a threshold reference for the convergence criteria.
Denote the incident angle of point $\mathbf{p}$ to the fitted plane $S$ as $\phi$, it can be expressed as $\cos^{-1}\tfrac{|\mathbf{u}\cdot\hat{\mathbf{p}}^{G}|}{|\mathbf{u}|\cdot|\hat{\mathbf{p}}^{G}|}$ according to Eq. 3. Note that $\mathbf{z}$ is the point-to-plane residual vector and $z$ is its scalar value. Assuming the ranging error of LiDAR point follows Gaussian distribution with zero mean and $\sigma$ variance. When the laser beam is perpendicular to the fitted plane, the residual $z_{\perp}$ follows chi-squared distribution and can be expressed as follows:
$\begin{gathered}{z}_{\perp}=|\mathbf{p}^{G}-\hat{\mathbf{p}}^{G}|\sim\mathcal{N}(0,\sigma)\\ \frac{{z}_{\perp}^{2}}{\sigma^{2}}\sim\chi^{2}(1),\;\mathbb{E}({z}_{\perp})=\sigma\end{gathered}$ (10)
As the surroundings vary during system’s ego-motion, the distribution of normal vector $\mathbf{u}$ to plane $S$ also changes. The variation of the geometrical scene would be reflected in the distribution of the incident angle $\phi$. From Fig. 3, the residual variance $\sigma_{z}$ should be $\sigma\cos\phi$. Assuming the incident angle is uniformly distributed over the domain $[0,\pi/2)$, then the expected mean of point-to-plane residual can be expressed as:
$\begin{gathered}\mathbb{E}({z})=\frac{2}{\pi}\int_{0}^{\frac{\pi}{2}}\mathbb{E}({z}_{\perp})\cos{\phi}\;d\phi=\frac{2}{\pi}\sigma\end{gathered}$ (11)
Therefore, assuming the pointcloud registration converges to the global optimum, and only considering the LiDAR ranging error, the expectation of APR should be ${2\sigma}/{\pi}$. If the mean value of current point-to-plane residual is lower than this criteria, the last registration can be regarded as sufficiently converged.
Based on this threshold of convergence criteria, we can easily determine the convergence degree of pointcloud registration, and control the backpropagation and asymptotic distortion compensation. Noting that $\sigma_{t}=\mathbb{E}({z})$ and the selected threshold of convergence criteria as $\eta$, due to the factors such as dynamic objects and cumulative errors in pointcloud map, the threshold $\eta$ can be appropriately extended to the interval of $[\sigma_{t},2\sigma_{t}]$ in practice.
In detail, we perform two conditional judgments in iterative state estimation: First, evaluate the previous LiDAR frame, if its APR is less than the threshold $\eta$, it can be considered to be sufficiently converged and we will continue. Then, check the convergence of current registration, if the APR is still less than $\eta$, there is no need to backpropagate for small residual distortion; otherwise, we will perform the backpropagation and compensation during the iteration.
On one hand, this approach can ensure the backpropagation and asymptotic compensation only performed as the previous LiDAR frame converged well, thus enhancing the odometry robustness in long-term and large-scale scenario. On the other hand, it can avoid lots of additional operations in unreliable or unnecessary conditions, to improve the system computational efficiency. The experiment results of our framework are shown in the next section.
IV EXPERIMENT RESULTS
TABLE II: Odometry Accuracy Evaluation - RMSE of ATE & End to End Errors($m$)
Dataset LiDAR Type Sequence Ours FAST-LIO2 DLIO SLICT Point-LIO
WHU1 Helmet Livox AVIA Residence_5.2 (452.9m) 0.415 0.505 4.660 1.550 $\times$
Park_3.2 (628.9m) 0.582 0.618 7.725 0.602 $\times$
Subway_3.3 (827.0m) 1.670 3.752 3.688 1.967 $\times$
Mall_6.3 (1006.5m) 0.216 0.208 3.238 0.274 15.409
Botanic1 Garden Livox AVIA 1005_05 (566.8m) 0.773 1.053 0.640 7.931 0.917
1006_03 (686.2m) 1.796 2.075 3.017 6.527 2.471
1008_01 (855.9m) 1.152 1.577 14.829 4.445 1.590
VLP16 1005_05 (566.8m) 0.175 0.244 0.565 37.425 0.750
1006_03 (686.2m) 0.377 0.388 0.393 $\times$ 0.445
1008_01 (855.9m) 0.339 0.363 0.694 $\times$ 0.457
Handheld2 Livox AVIA Indoor_1 (~100m) 0.893 1.133 $\times$ 1.960 5.401
Indoor_2 (~100m) 0.227 0.204 3.354 0.249 0.164
Indoor_3 (~100m) 1.596 3.001 $\times$ 2.476 6.844
Outdoor_1 (~300m) 0.138 1.901 0.486 1.835 0.728
Outdoor_2 (~400m) 0.984 2.194 7.879 1.803 1.261
Outdoor_3 (~500m) 2.666 4.702 1.445 4.127 4.607
Average(RMSE) 0.750 1.078 3.945 7.590 3.148
  • 1  means the dataset is evaluated by RMSE of ATE, and  2 means it is evaluated by End to End errors.
  • $\times$  denotes the system severely diverged midway, and $\underline{\hskip 7.11317pt}$  denotes the second best result compared to bolded result.
In this section, we qualitatively and quantitatively evaluate AC-LIO framework through a series of experiments with various public and self-collected datasets. To examine the performance of our approach in long-term and large-scale scenarios, we conduct extensive experiments on the odometry accuracy, the time efficiency of system, and the validity of convergence criteria.
Figure 4: Our handheld sensor suite used in self-collected dataset, including the solid-state LiDAR Livox Avia with the built-in BMI088 IMU, and three other cameras.
IV-A Experiment Dataset
The datasets cover a wide range of structured environments (e.g., campus road and garage) and unstructured settings (e.g., botanical gardens), as well as different carriage modes (helmet, vehicle, handheld) to boost the challenge and completeness of assessment.
  • WHU-Helmet [40]: Collected on compact helmet system with Livox LiDAR, the dataset contains complex large-scale scenes including wild, urban and tunnel, with challenges of drastic motion from wearer and repeated symmetrical geometric structures.
  • BotanicGarden [41]: Mounted on all-terrain wheeled robot, this dataset consists of diverse natural scenes including riversides, narrow trails and bridges with both Livox Avia and VLP16 in sequence.
  • Handheld datasets: We collected multiple datasets of different scenarios and motion conditions, covering indoor and outdoor scenes like library, campus road and UAV airfield with Livox Avia. The handheld data collection device is shown in Fig. 4.
IV-B Odometry Accuracy Evaluation
Figure 5: Accuracy evaluation in campus road scene. The figure shows the result of sequence Outdoor_1. Within the dashed box are, in order, the car, the streetlight, and the pattern on the ground. We keep running after initialization to enhance the motion intensity and complete a closed loop about 300m. The closure trajectory and the map consistency suggest that our AC-LIO achieves better performance.
In odometry accuracy experiments, we compare AC-LIO framework with current state-of-the-art LIO systems including FAST-LIO2, DLIO, SLICT and Point-LIO. Some other frameworks are designed for spinning LiDAR hence they’re not included in comparison. All algorithms are implemented in C++ and employed on an i9 CPU computer with the ROS in Ubuntu. These methods only utilize CPU for computation without the assistance of GPU parallel computing. For comparison, we try to fine-tune other methods’ parameters to achieve their best performance. Besides, we maintain the threshold of convergence criteria in the extended interval, i.e., $\eta\in[\sigma_{t},2\sigma_{t}]$.
In the evaluation of WHU-Helmet, the sequences Residence_5.2 and Subway_3.3 introduce aggressive motion combined with locally degraded scenarios, while Park_3.2 and Mall_6.3 suffer from extensive occlusion by pedestrians. These factors impose higher demands on the robustness of the LIO system. For the BotanicGarden datasets, we test both the solid-state (Livox AVIA) and the spinning (VLP16) LiDAR. Since the sensors are mounted on a wheeled robotic platform, the overall motion is smooth and suitable for evaluating long-duration drift under regular motion conditions. However, the abundance of unstructured scenes in the garden (e.g., thick woods, bushes) poses challenges for constructing residual constraints. In our self-collected datasets, we handheld the sensor suite and completed multiple closed-loop paths in different scenarios. In Indoor_1 UAV airfield and Indoor_3 library scenarios, we conduct intense turning in narrow environment. For the Outdoor sequences, we handhold the device and attempt different motion conditions including running and sharp turns in campus road scenes.
Concrete experiment results are demonstrated in Table II and Fig. 5. Our approach further optimizes the residual motion distortion via feedback mechanism based on the convergence criteria, thus improving the accuracy while ensuring the robustness of the system under challenging scenarios. The results indicate that our AC-LIO method generally outperforms other benchmark frameworks in terms of accuracy during long-term, large-scale scenes, with the average RMSE among sequences decreased by 30.4% compared to the second best, and exhibits high robustness under challenging scenarios.
TABLE III: Time Consumption Evaluation ($ms$)
Ours FAST-LIO2 DLIO SLICT Point-LIO
Min 4.65 3.15 3.43 21.33 11.95
Max 74.46 28.48 27.01 345.97 62.36
Mean 14.62 16.09 9.12 61.97 34.31
Median 11.71 16.68 8.50 51.49 34.98
Figure 6: The result of time efficiency evaluation. The cross line inside box denotes the median process time. Our AC-LIO achieves the second highest efficiency, only behind the lightweight DLIO framework. And the maximum iteration time also meets the real-time requirements.
IV-C Time Efficiency Evaluation
We calculate the runtime of each algorithm during iteration, with the results shown in Table III and Fig. 6. In terms of average and median processing time, the time efficiency of our AC-LIO framework ranks second place among the tested frameworks, just following the lightweight DLIO framework, and the maximum time for iteration also satisfies realtime requirement (below 100ms for 10Hz LiDAR input). Due to the convergence criteria strategy, our method can avoid incurring extra optimization costs in unreliable or unnecessary conditions, thus achieving the balance between system convergence and high time efficiency.
IV-D Convergence Criteria Validity Evaluation
TABLE IV: Convergence Criteria Validity Evaluation - RMSE of ATE ($m$)
00 0 $0.5\sigma_{t}$ $[\sigma_{t},2\sigma_{t}]$ $3\sigma_{t}$
Park_3.2
0.624 0.610 0.582 0.616
Subway_3.3
2.841 2.419 1.670 1.788
Garden_1006
1.901 1.709 1.796 1.892
Garden_1008
1.197 1.199 1.152 1.178
To evaluate the validity of the proposed convergence criteria, we examine the odometry accuracy at different $\eta$ threshold levels. While keeping other parameters the same, we only adjust the threshold of convergence criteria. Due to the factors such as dynamic objects, cumulative pointcloud map errors, etc., we fine-tune the threshold $\eta$ within the interval of $[\sigma_{t},2\sigma_{t}]$. The specific experimental results are shown in Table IV.
  • $\eta=0$: No backpropagation and asymptotic compensation is performed at this situation, reflecting the basic accuracy of the odometry.
  • $\eta=0.5\sigma_{t}$: Lower than the expected mean of pointcloud residuals, rarely perform backpropagation. (Only considering LiDAR ranging errors, $\mathbb{E}({z})=\sigma_{t}$)
  • $\eta\in[\sigma_{t},2\sigma_{t}]$: Recommended interval in consideration of other error factors, assuring the backpropagation and asymptotic compensation will be performed with sufficient convergence in previous frame.
  • $\eta=3\sigma_{t}$: Too high a threshold may lead to backpropagation in case of large initial errors, potentially damaging the system’s accuracy and robustness.
The experimental results show that, in most cases, selecting the threshold $\eta$ within the range of $[\sigma_{t},2\sigma_{t}]$ achieved higher accuracy. Overall, as the threshold increases from zero, the odometry accuracy initially improves because more sufficiently converged states are allowed to perform backpropagation and asymptotic compensation. However, as the threshold increases further, states with larger initial errors are also allowed to backpropagate, which reduces the effectiveness of the asymptotic compensation and negatively impacts odometry accuracy.
V Conclusion and Future Work
In this paper, we proposed AC-LIO, a novel asymptotically and consistently converging LIO framework. We reconsider the factors that may introduce errors in long-duration, large-scale scenarios, aiming to further enhance the convergence of LIO system. The approach of guiding asymptotic compensation based on convergence criteria is designed to promote consistent convergence of pointcloud registration, further improving the odometry accuracy while ensuring system robustness. Numerous experiment results indicate that our framework’s average RMSE is reduced by about 30.4% over the second best result, achieving higher accuracy in long-term, large-scale localization and mapping.
Future Work: We notice the performance of back propagation depends on the guidance of convergence criteria to some extend. The introduction of convergence criteria is essentially an attempt to overcome the disruption of the initial error on the back propagation during the iteration update. In future work, we will continue to study the convergence assessment of odometry, the combination of asymptotic backpropagation with other strategies, and the perception and adaptation of degenerate scenarios, to further explore the error-analyzable and scenario-interpretable SLAM framework.